OpenStudy (anonymous):

Cameron is replacing the roof on a gazebo. The roof is in the shape of a cone with an open bottom. The diameter is fourteen feet and the slant height is ten feet. What is the lateral area of the roof? Use three point one four for pi.

OpenStudy (e.cociuba):

lateral area = pi*r*s where s is the length of the slope of the cone pi*r*s = pi*10*14 =140 *pi = ___?___ ft^2

OpenStudy (e.cociuba):

So you do 140 x 3.14(pi) and you get 439.6 ft^2

OpenStudy (anonymous):

OH. At first I was like *.* then I got it. Thank you~ c':

OpenStudy (e.cociuba):

Are you sure your positive that you understand? I will explain if needed :)

OpenStudy (anonymous):

I am sure ^-^ Thank you~ I will be sure to use this for future references

OpenStudy (e.cociuba):

No Problemo! :)